Add the following lengths. Give the answer in feet and inches in its simplest form. (I am having trouble converting the mixed numbers)
6ft.7in. 47in. 9ft.11in. 13 3/4ft. 18ft.
9 2/3ft. 17in.

Example:
6' 7''
0' 47''
9' 11''

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

6FT7IN
0FT47IN
9FT11IN
13FT9IN
18FT0IN
9FT8IN
0FT17IN
------------------------------------
55FT99IN 99/12=8FT 3IN.
55FT
8FT3IN
-------------------------------
63FT3IN. ANSWER.

Answer by stanbon(75887)

Give the answer in feet and inches in its simplest form. (I am having trouble converting the mixed numbers)
6ft.7in. 47in. 9ft.11in. 13 3/4ft. 18ft.
9 2/3ft. 17in.
---------------
Convert each one to inches;
Add the inches:
Divide by 12 to get the number of feet plus some number of inches.
=========================================
6'7" = 79 inches
47" = 47 inches
9'11" =119 inches
[13(3/4)]' =165 inches
18' =216 inches
[9 2/3]' =116 inches
17' = 17 inches
=====================
Add to get: 759"
= 63 ft 3 inches
==================
Cheers,
Stan H.